A firm has developed forecasts for the yen for each of the four quarters of a...

  1. A firm has developed forecasts for the yen for each of the four quarters of a year.

Quarter      Yen forecast        Yen actual value     Pound forecast    Pound actual value

1              $.0050                     $.0051                  $1.50                      $1.51

2              $.0048                     $.0052                  $1.53                      $1.50  

3              $.0053                     $.0052                  $1.55                      $1.58

  1.           $.0055                     $.0056                  $1.49                      $1.52
  1. Calculate the absolute forecast errors in percentage terms for each quarter for both the yen and the pound. Then calculate the mean absolute forecast error over the four years for each currency.

  1. Which currency forecasts were more accurate?
Answer #1

in % 1. fore cast errors terms = Cal Absolute actual - for cast x 100 Actual Quarter yen Yen forecast Actual value I absoluteFor pound = 0.664 + 2), t l90%. for pound 4.56%. = 1,524, (b) Pound Currencys for cast were more accurate its lower mean fo